OpenStudy (anonymous):

How do I state the slope and y-intercept in this equation 3x + 5y -30 = 0? Please help

OpenStudy (shamil98):

Put the equation in slope intercept form, y= mx + b

OpenStudy (tester97):

first subtract 3x from eash side then what are you left with?

OpenStudy (shamil98):

3x + 5y - 30 = 3x + 5y = 30 5y = -3x + 30 y = -3/5x + 6 The y intercept is 6 and the slope is -3/5

OpenStudy (anonymous):

Wow thanx

OpenStudy (anonymous):

Y subtract 3x?

OpenStudy (tester97):

because you are trying to change it into slope intercept form then you would have 5y=-3x-30

OpenStudy (tester97):

then you would divide by 5 on both sides so you could have y by itself on one side so you would have y= -3/5x-6

OpenStudy (anonymous):

Thanx alot I got ut now

OpenStudy (anonymous):

It* sorry

OpenStudy (anonymous):

I just put another question can u guys help me plz?